OpenStudy (anonymous):

calculate the divergence of vector field v(x,y,z)= (e^x , ln(xy) , e^(xyz)) = e^x i + ln(xy) j + e^(xyz)k and evaluate it at the point P = (2,1,-1) div v = ? div v (2,1,-1) = ?

OpenStudy (amistre64):

http://www.youtube.com/watch?v=AJsPwNgJoKI this guy tends to do a good demonstration

OpenStudy (amistre64):

Vx = e^x Vy = 1/y Vz = xy e^(xyz) e^x , 1/y , xy e^(xyz) e^x , ln(xy) , e^(xyz) ----------------------- e^(2x) + ln(xy)/y +xy e^(2xyz), is the divergence then

OpenStudy (amistre64):

assuming i did my partials correctly

OpenStudy (amistre64):

the rest is just plugging in x,y,z from the point given

OpenStudy (anonymous):

good video that explains it well! thank you very much

OpenStudy (amistre64):

yeah, i like that guy; and also patrick on youtube; very good instructors